HOW much water must be mixed with 9 pints of 70% developer to produce a mixture that is 25% developer?
Answer by checkley77(12569)

.70*9=.25(9+x)
6.3=2.25+.25x
.25x=6.3-2.25
.25x=4.05
x=4.05/.25
x=16.2 pints of water is needed.
Proof:
.70*9=.25(9+16.2)
6.3=.25*25.2
6.3=6.3